Comparative data on three companies in the same service industry are given below Required: 2. Fill in the missing information. (Round the Turnover, ROI answers to 2 decimal places.) Companies in the Same Industry $ 3,264,000 $ 522,240 $ 1,356,000 $ 244,080 Sales Net operating income Average operating assets$1,920,000 Margin Turnover Return on investment (ROI) $ 2,890,000 41% 2.10 7.20
Answer #1
Companies in the Same Industry
A B C
Sales $ 3,264,000 $ 1,356,000 $ 6,069,000
Net operating income $ 522,240 $ 244,080 $ 242,760
Average operating assets $ 1,920,000 $ 3,390,000 $ 2,890,000
Margin 16 % 18 % 4 %
Turnover 1.70 0.40 2.10
Return on investment (ROI) 27.20 % 7.20 % 8.40 %

Working Notes:

Company A

Margin=Net operating income/Sales*100

Margin=$ 522,240/$ 3,264,000*100=16%

Turnover=Sales/Average operating assets

Turnover=$ 3,264,000/$ 1,920,000=1.70

Return on investment (ROI)=Net operating income/Average operating assets

ROI=$ 522,240/$ 1,920,000=27.20%

Company B

Average operating assets=Net operating income/ROI

Average operating assets=$ 244,080/7.20%=$ 3,390,000

Margin=Net operating income/Sales*100

Margin=$ 244,280/$ 31,356,000*100=18%

Turnover=Sales/Average operating assets

Turnover=$ 1,356,000/$ 3,390,000=0.40

Company C

Sales=Turnover*Average operating assets

Sales=2.10*$ 2,890,000=$ 6,069,000

Net operating income=Sales*Margin

Net operating income=$ 6,069,000*4%=$ 242,760

Return on investment (ROI)=Net operating income/Average operating assets

ROI=$ 242,760/$ 2,890,000=8.40%